# Fast Sign Retrieval via Sub-band Convolution: An Elementary Extension of Binary Classification

### v2

@@ 1 Introduction
Various sign retrieval methods have been proposed, as summarized in Fig. 2(a). To the best of our knowledge, all of the previous methods, without exception, are based on iterative optimization. For example, in [7, 8], an image and signed DCT coefficients are alternatively retrieved, and the signs of the retrieved coefficients are extracted. In [14], the signs are regarded as variables and are optimized in an iterative manner. Previous methods have demonstrated the ability to accurately retrieve a large portion of the signs; for example, as reported in [8], roughly $75$% of the total signs can be retrieved. However, these methods come at an excessive computation cost, primarily due to computing the DCT and the inverse DCT (IDCT) for each iteration. For example, as reported in [8], it takes approximately 10 [s] to retrieve $256\times 256$ signs, making sign retrieval impractical for real-time encoding and decoding scenarios.
@@ 3.2 Sign retrieval via sub-band convolution
where $u(z)=\lfloor z/8\rfloor$, $v(z)=z\,\text{mod}\,8$, $m(x)=x$, and $n(y)=y$. The amplitudes/signs belonging to the $(u(z),v(z))$-th frequency band are packed in the $z$-th 2D slice, referred to as the sub-band block, and the 3D tensors contain all the $64$ sub-band blocks, as shown in Figure 3. We simply stack these $64$ sub-bands along the $z$ dimension; consideration of even/odd symmetry of the DCT kernels might lead to a better method for handling these sub-bands, but we leave it as the future work. The effectiveness of the reshaping process will be demonstrated in Section 4.2.
@@
At the test phase, we obtain the retrieved signs by thresholding $\mathsf{F}_{x,y,z}(\mathsf{A},\Theta^{*})$ as follows.

### v4

@@ 1 Introduction
Various sign retrieval methods have been proposed, as summarized in Fig. 2(a). To the best of our knowledge, all of the previous methods, without exception, are based on iterative optimization. For example, in [7, 8], an image and signed DCT coefficients are alternatively retrieved, and the signs of the retrieved coefficients are extracted. In [14], the signs are regarded as variables and are optimized in an iterative manner. Previous methods have demonstrated the ability to accurately retrieve a large portion of the signs; for example, as reported in [8], roughly $75$% of the total signs can be retrieved. However, these methods come at an excessive computation cost, primarily due to computing the DCT and the inverse DCT (IDCT) for each iteration. For example, as reported in [8], it takes approximately 10 s to retrieve $256\times 256$ signs, making sign retrieval impractical for real-time encoding and decoding scenarios.
@@ 3.2 Sign retrieval via sub-band convolution
where $u(z)=\lfloor z/8\rfloor$, $v(z)=z\,\text{mod}\,8$, $m(x)=x$, and $n(y)=y$. The amplitudes/signs belonging to the $(u(z),v(z))$-th frequency band are packed in the $z$-th 2D slice, referred to as the sub-band block, and the 3D tensors contain all the $64$ sub-band blocks, as shown in Fig. 3. We simply stack these $64$ sub-bands along the $z$ dimension; consideration of even/odd symmetry of the DCT kernels might lead to a better method for handling these sub-bands, but we leave it as the future work. The effectiveness of the reshaping process will be demonstrated in Section 4.2.
@@
At the test phase, we obtain the retrieved signs by thresholding $\mathsf{F}_{x,y,z}(\mathsf{A},\tilde{\Theta})$ as follows.